OpenStudy (anonymous):

find the values when n is equal or greater than 1 for which 1!+2!+3!+...+n! is a perfect square

OpenStudy (anonymous):

\[1!+2!+3!\] works ,you get 9

OpenStudy (anonymous):

yes but i am looking for all the values

OpenStudy (raden):

n=1 :)

OpenStudy (raden):

i think, Only 1 and 3. If n =>4 the unit digit of the expression is always 3 . So u can conclude that there are no perfect square of the form 1!+2!+...+n! If n=4,5,6,....

OpenStudy (anonymous):

\[ n! + (n+1)! = n!+(n+1)n! = (n+2)n! \]

OpenStudy (anonymous):

\[ (n+2)n! + (n+2)! = (n+2)n!+(n+2)(n+1)n! = (n+2)(n+2)^2n! \]Interesting pattern.

OpenStudy (anonymous):

Should be\[ (n+2)n! + (n+2)! = (n+2)n!+(n+2)(n+1)n! = (n+2)^2n! \]

OpenStudy (raden):

so, ?

OpenStudy (anonymous):

I actually don't understand your proof.

OpenStudy (anonymous):

You're saying last digit must be 3?

OpenStudy (raden):

1! = 1 = 1^2 (true it is a perfect square) 1! + 2! = 3 is not a perfect square 1! + 2! + 3! = 9 = 3^2 (true) 1! + 2! + 3! + 4! = 23 is not a perfect square 1! + 2! + 3! + 4! + 5! = 143 is not a perfect square 1! + 2! + 3! + 4! + 5! + 6! = 863 is not a perfect square ..... and so on, will has the unit always 3 while, no square numbers has a unit "3" 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = ...6 5^2 = ...5 6^2 = ...6 7^2 = ...9 8^2 = ...4 9^2 = ...1 10^2 = ... 0

OpenStudy (anonymous):

1 and 3

OpenStudy (anonymous):

agree with @RadEn :) there is no positive integer m such that\[m^2 \equiv 3 \ \ \ (\text{mod} \ 10) \]on the other hand for \(n\ge 5\)\[1!+2!+3!+...+n! \equiv \ 1!+2!+3!+4!\equiv33 \equiv 3 \ \ (\text{mod} 10)\]trying \(n=1,2,3,4\) gives \(n=1,3\) as answer.